# SCAM: A Real-World Typographic Robustness Evaluation for Multimodal Foundation Models

### v1

@@ Abstract
Typographic attacks exploit the interplay between text and visual content in multimodal foundation models, causing misclassifications when misleading text is embedded within images. However, existing datasets are limited in size and diversity, making it difficult to study such vulnerabilities. In this paper, we introduce SCAM, the largest and most diverse dataset of real-world typographic attack images to date, containing 1,162 images across hundreds of object categories and attack words. Through extensive benchmarking of Vision-Language Models (VLMs) on SCAM, we demonstrate that typographic attacks significantly degrade performance, and identify that training data and model architecture influence the susceptibility to these attacks. Our findings reveal that typographic attacks persist in state-of-the-art Large Vision-Language Models (LVLMs) due to the choice of their vision encoder, though larger Large Language Models (LLMs) backbones help mitigate their vulnerability. Additionally, we demonstrate that synthetic attacks closely resemble real-world (handwritten) attacks, validating their use in research. Our work provides a comprehensive resource and empirical insights to facilitate future research toward robust and trustworthy multimodal AI systems. We publicly release the datasets introduced in this paper under https://huggingface.co/datasets/BLISS-e-V/SCAM, along with the code for evaluations at https://github.com/Bliss-e-V/SCAM.
@@ 1 Introduction
RN50 openai 97.50 36.32 ↓61.18
ViT-B-32 laion2b 97.68 74.18 ↓23.50
ViT-B-16 laion2b 98.54 68.93 ↓29.61
ViT-B-16-SigLIP webli 99.05 81.50 ↓17.55
ViT-L-14 commonpool_xl 99.40 74.96 ↓24.44
openai 98.80 41.05 ↓57.75
ViT-L-14-336 openai 99.23 34.68 ↓64.55
ViT-g-14 laion2b 98.71 62.05 ↓36.66
ViT-bigG-14 laion2b 99.40 70.91 ↓28.49
llava-llama3-1.1:8b - 97.83 40.17 ↓57.66
llava-1.6:7b - 96.30 57.78 ↓38.52
llava-1.6:13b - 97.86 62.42 ↓35.44
llava-1.6:34b - 98.62 86.67 ↓11.95
llama3.2-vision:90b - 99.05 70.75 ↓28.30
gpt-4o-mini-2024-07-18 - 99.14 91.39 ↓7.75
gpt-4o-2024-08-06 - 99.57 97.41 ↓2.16
@@ 3 SCAM Dataset
To create the cleaned NoSCAM dataset, we remove all attack words by obscuring the post-it notes while preserving the original object images. Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font) and matching the colors of the post-its and ink from the original images.
@@ 5 Results
As shown in fig. 4, VLMs achieve high prediction accuracy in the zero-shot classification task on the NoSCAM dataset. However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets. Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 35 percentage points. For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
@@
commonpool_l_text 91.57
commonpool_l_image 90.71
commonpool_l_basic 89.76
commonpool_l 88.47
datacomp_l 79.60
commonpool_l_laion 78.83
commonpool_l_clip 70.91
laion2b 68.93
laion400m 68.80
datacomp_xl 68.07
@@
openai CLIP-ViT 44.06
CLIP-RN 35.92
v1 SigLIP-ViT 75.77
CLIP-ViT 69.02
@@
Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table 1), the strength of their vision encoder plays a significant role in this susceptibility. Smaller LLaVA models, such as llava-llama3:8b, llava-1.6:7b, and llava-1.6:13b, suffer substantial accuracy drops of more than 30-50 percentage points under typographic attacks. This vulnerability largely stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig. 6). However, as model size increases, the robustness to typographic attacks improves. LLaVA models with larger LLM backbones, such as llava-1.6:34b, exhibit significantly better performance under attack. This suggests that a more powerful LLM backbone can compensate for the limitations of the vision encoder, making these models more resilient to typographic distortions. In comparison, models trained by OpenAI, such as gpt-4o, exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 2 percentage points. The smaller gpt-4o-mini, however, experiences a higher drop of approximately 8 points, illustrating again that larger models tend to be more robust. Despite that, model size alone does not determine robustness. In fact, the large llama3.2-vision:90b model is more vulnerable to typographic attacks than the smaller llava-1.6:34b, highlighting the critical role of the vision encoder in overall performance. These results highlight that while vision encoders remain a critical vulnerability, increasing the size and quality of the LLM backbone plays a key role in reducing susceptibility to typographic attacks.
@@ 7 Conclusion
JW was responsible for image processing, code development, model evaluations, visualizations, and overall coordination. EP contributed to conceptualization, figure design, and manuscript writing, including the introduction, dataset, evaluation, and drafted most of the results and discussion sections. JH focused on related work research and supported overall writing. LP contributed to dataset expansion through large-scale labeling and provided continuous critical feedback. LH initiated the project, provided supervision throughout, designed the data collection protocol, and guided the conceptual and methodological direction of the work. All authors contributed to image collection and labeling.
@@ A.4 Model Accuracies
Table 9: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM datasets.
@@
NoSCAM SCAM SynthSCAM
RN101 openai 97.42 40.02 27.71
RN50 openai 97.50 36.32 24.27
RN50x16 openai 98.62 37.69 35.28
RN50x4 openai 97.85 33.05 29.69
RN50x64 openai 99.05 32.53 32.70
ViT-B-16 commonpool_l_basic_s1b_b8k 98.54 89.76 82.87
commonpool_l_clip_s1b_b8k 98.11 70.91 63.08
commonpool_l_image_s1b_b8k 98.19 90.71 86.14
commonpool_l_laion_s1b_b8k 97.42 78.83 67.30
commonpool_l_s1b_b8k 97.50 88.47 81.33
commonpool_l_text_s1b_b8k 98.02 91.57 85.97
datacomp_l_s1b_b8k 98.02 79.60 69.62
datacomp_xl_s13b_b90k 98.80 68.07 55.34
laion2b_s34b_b88k 98.54 68.93 52.75
laion400m_e31 98.02 68.93 52.93
laion400m_e32 98.11 68.67 53.10
metaclip_400m 97.85 58.78 37.95
metaclip_fullcc 98.71 61.10 42.17
openai 97.07 40.02 22.03
ViT-B-16-SigLIP webli 99.05 81.50 76.25
ViT-B-16-SigLIP-256 webli 99.05 81.07 75.90
ViT-B-16-SigLIP-384 webli 99.40 84.42 80.98
ViT-B-16-SigLIP-512 webli 99.57 85.71 81.93
ViT-B-16-SigLIP-i18n-256 webli 99.48 89.41 86.14
ViT-B-16-plus-240 laion400m_e31 98.45 71.26 55.08
laion400m_e32 98.45 71.26 55.08
ViT-B-32 commonpool_m_basic_s128m_b4k 89.16 90.71 91.48
commonpool_m_clip_s128m_b4k 90.19 90.62 90.53
commonpool_m_image_s128m_b4k 89.07 91.31 92.69
commonpool_m_laion_s128m_b4k 87.26 90.10 90.02
commonpool_m_s128m_b4k 87.01 89.67 91.31
commonpool_m_text_s128m_b4k 88.55 90.96 91.22
commonpool_s_basic_s13m_b4k 72.03 74.27 74.96
commonpool_s_clip_s13m_b4k 76.51 76.42 76.33
commonpool_s_image_s13m_b4k 72.29 73.06 73.41
commonpool_s_laion_s13m_b4k 71.69 73.49 74.01
commonpool_s_s13m_b4k 73.32 74.01 73.84
commonpool_s_text_s13m_b4k 73.49 74.01 74.61
datacomp_m_s128m_b4k 89.16 90.62 91.05
datacomp_xl_s13b_b90k 98.11 78.66 65.40
laion2b_s34b_b79k 97.68 74.18 58.18
laion400m_e31 95.78 63.34 42.43
laion400m_e32 95.70 63.25 42.69
metaclip_400m 96.82 76.59 59.21
metaclip_fullcc 97.59 79.43 61.53
openai 96.82 60.50 35.89
ViT-B-32-256 datacomp_s34b_b86k 98.11 73.15 60.93
ViT-H-14 laion2b_s32b_b79k 98.88 63.08 49.66
metaclip_altogether 99.40 66.87 54.48
metaclip_fullcc 99.31 62.65 44.66
ViT-H-14-CLIPA datacomp1b 99.48 65.06 58.18
ViT-H-14-CLIPA-336 datacomp1b 99.57 61.62 56.54
laion2b 99.31 72.03 60.41
ViT-L-14 commonpool_xl_clip_s13b_b90k 99.05 57.75 50.09
commonpool_xl_laion_s13b_b90k 99.23 74.96 61.10
commonpool_xl_s13b_b90k 99.40 74.96 69.54
datacomp_xl_s13b_b90k 99.48 60.76 54.04
laion2b_s32b_b82k 98.80 62.91 50.77
laion400m_e31 98.97 67.38 53.01
laion400m_e32 98.97 66.95 53.36
metaclip_400m 98.45 58.00 41.65
metaclip_fullcc 99.14 63.43 46.39
openai 98.80 41.05 29.00
ViT-L-14-336 openai 99.23 34.68 27.45
ViT-L-14-CLIPA datacomp1b 99.40 72.81 64.63
ViT-L-14-CLIPA-336 datacomp1b 99.57 74.87 65.92
ViT-L-16-SigLIP-256 webli 99.40 86.75 81.24
ViT-L-16-SigLIP-384 webli 99.57 87.26 82.79
ViT-SO400M-14-SigLIP webli 99.57 84.51 83.30
ViT-SO400M-14-SigLIP-378 webli 99.66 85.80 84.94
ViT-SO400M-14-SigLIP-384 webli 99.66 85.63 84.94
ViT-SO400M-16-SigLIP-i18n-256 webli 99.40 88.81 84.17
ViT-bigG-14 laion2b_s39b_b160k 99.40 70.91 59.47
metaclip_fullcc 99.14 68.42 55.51
ViT-bigG-14-CLIPA datacomp1b 99.48 68.16 58.95
ViT-bigG-14-CLIPA-336 datacomp1b 99.57 64.03 57.14
ViT-g-14 laion2b_s12b_b42k 98.62 69.36 57.49
laion2b_s34b_b88k 98.71 62.05 48.02
convnext_base laion400m_s13b_b51k 98.11 58.69 41.22
convnext_base_w laion2b_s13b_b82k 98.71 66.52 55.08
laion2b_s13b_b82k_augreg 98.62 68.42 58.09
laion_aesthetic_s13b_b82k 98.11 67.56 54.48
convnext_base_w_320 laion_aesthetic_s13b_b82k 98.45 68.67 52.67
laion_aesthetic_s13b_b82k_augreg 98.36 71.94 61.27
convnext_large_d laion2b_s26b_b102k_augreg 99.23 67.30 56.02
convnext_large_d_320 laion2b_s29b_b131k_ft 99.05 67.04 58.09
laion2b_s29b_b131k_ft_soup 99.23 67.47 56.88
convnext_xxlarge laion2b_s34b_b82k_augreg 99.23 65.06 56.28
laion2b_s34b_b82k_augreg_rewind 99.05 66.27 55.34
laion2b_s34b_b82k_augreg_soup 99.23 65.15 55.85
nllb-clip-base v1 93.80 76.33 62.91
nllb-clip-base-siglip mrl 98.11 71.77 65.32
v1 98.36 77.54 76.68
nllb-clip-large v1 96.90 61.70 50.60
nllb-clip-large-siglip mrl 98.62 72.38 68.33
v1 99.05 74.01 76.16
roberta-ViT-B-32 laion2b_s12b_b32k 96.90 75.04 58.78
xlm-roberta-base-ViT-B-32 laion5b_s13b_b90k 97.85 76.25 60.76
xlm-roberta-large-ViT-H-14 frozen_laion5b_s13b_b90k 98.80 63.77 50.00
llava-llama3:8b - 97.83 40.17 43.88
llava:7b - 96.30 57.78 62.85
llava:13b - 97.68 62.42 56.66
llava:34b - 98.62 86.67 85.55
llama3.2-vision:90b - 99.05 70.75 72.33
gpt-4o-mini-2024-07-18 - 99.14 91.39 91.82
gpt-4o-2024-08-06 - 99.57 97.41 96.82

### v3

@@ Abstract
Typographic attacks exploit the interplay between text and visual content in multimodal foundation models, causing misclassifications when misleading text is embedded within images. However, existing datasets are limited in size and diversity, making it difficult to study such vulnerabilities. In this paper, we introduce SCAM, the largest and most diverse dataset of real-world typographic attack images to date, containing 1,162 images across hundreds of object categories and attack words. Through extensive benchmarking of Vision-Language Models (VLMs) on SCAM, we demonstrate that typographic attacks significantly degrade performance, and identify that training data and model architecture influence the susceptibility to these attacks. Our findings reveal that typographic attacks persist in state-of-the-art Large Vision-Language Models (LVLMs) due to the choice of their vision encoder, though larger Large Language Models (LLMs) backbones help mitigate their vulnerability. Additionally, we demonstrate that synthetic attacks closely resemble real-world (handwritten) attacks, validating their use in research. Our work provides a comprehensive resource and empirical insights to facilitate future research toward robust and trustworthy multimodal AI systems. We publicly release the datasets introduced in this paper along with the code for evaluations at www.bliss.berlin/research/scam.
@@ 1 Introduction
RN50 openai 97.76 36.61 ↓61.15
ViT-B-32 laion2b 98.45 74.68 ↓23.77
ViT-B-16 laion2b 98.71 69.16 ↓29.55
ViT-B-16-SigLIP webli 99.22 81.40 ↓17.82
ViT-L-14 commonpool_xl 99.48 74.68 ↓24.80
openai 99.14 40.14 ↓59.00
ViT-L-14-336 openai 99.22 33.85 ↓66.37
ViT-g-14 laion2b 99.05 61.93 ↓37.12
ViT-bigG-14 laion2b 99.40 70.89 ↓28.51
llava-llama3-1.1:8b - 98.09 39.50 ↓58.59
llava-1.6:7b - 97.50 58.43 ↓39.07
llava-1.6:13b - 98.88 58.00 ↓40.88
llava-1.6:34b - 98.97 84.85 ↓14.11
llama3.2-vision:90b - 98.88 71.01 ↓27.87
gpt-4o-mini-2024-07-18 - 99.40 84.68 ↓14.72
gpt-4o-2024-08-06 - 99.48 96.82 02.67
@@ 3 SCAM Dataset
To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images. Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images.
@@ 5 Results
As shown in fig. 4, VLMs achieve high prediction accuracy in the zero-shot classification task on the NoSCAM dataset. However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets. Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 36 percentage points. For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
@@
commonpool_l_text 91.82
commonpool_l_image 90.96
commonpool_l_basic 90.09
commonpool_l 88.72
datacomp_l 80.10
commonpool_l_laion 79.24
commonpool_l_clip 71.15
laion2b 69.16
laion400m 69.04
datacomp_xl 68.39
@@
openai CLIP-ViT 43.63
CLIP-RN 35.76
v1 SigLIP-ViT 75.80
CLIP-ViT 69.04
@@
Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table 1), the strength of their vision encoder plays a significant role in this susceptibility. Smaller LLaVA models, such as llava-llama3:8b, llava-1.6:7b, and llava-1.6:13b, suffer substantial accuracy drops of more than 40-60 percentage points under typographic attacks. This vulnerability largely stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig. 6). However, as model size increases, the robustness to typographic attacks improves. LLaVA models with larger LLM backbones, such as llava-1.6:34b, exhibit significantly better performance under attack. This suggests that a more powerful LLM backbone can compensate for the limitations of the vision encoder, making these models more resilient to typographic distortions. In comparison, models trained by OpenAI, such as gpt-4o, exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 3 percentage points. The smaller gpt-4o-mini, however, experiences a higher drop of approximately 15 points, illustrating again that larger models tend to be more robust. Despite that, model size alone does not determine robustness. In fact, the large llama3.2-vision:90b model is more vulnerable to typographic attacks than the smaller llava-1.6:34b, highlighting the critical role of the vision encoder in overall performance. These results highlight that while vision encoders remain a critical vulnerability, increasing the size and quality of the LLM backbone plays a key role in reducing susceptibility to typographic attacks.
@@ 7 Conclusion
JW was responsible for image processing, code development, model evaluations, visualizations, and overall coordination. EP contributed to conceptualization, figure design, and manuscript writing, including the introduction, dataset, evaluation, and drafted most of the results and discussion sections. JH focused on related work research and supported overall writing. JL supported model evaluations, managed data and code hosting, and created the project webpage. LP contributed to dataset expansion through large-scale labeling and provided continuous critical feedback. LH initiated the project, provided supervision throughout, designed the data collection protocol, and guided the conceptual and methodological direction of the work. All authors contributed to image collection and labeling.
@@ A.4 Model Accuracies
Table 9: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM, PAINT [38], and RTA-100 [36] datasets.
@@
NoSCAM SCAM SynthSCAM PAINT RTA-100
RN101 openai 97.93 40.14 26.18 36.36 35.10
RN50 openai 97.76 36.61 22.39 33.64 30.40
RN50x16 openai 98.79 37.12 34.19 46.36 32.70
RN50x4 openai 98.19 32.90 26.36 35.45 30.50
RN50x64 openai 98.97 32.04 31.78 43.64 35.40
ViT-B-16 commonpool_l_basic_s1b_b8k 98.62 90.09 83.03 82.73 88.10
commonpool_l_clip_s1b_b8k 98.54 71.15 62.53 69.09 71.50
commonpool_l_image_s1b_b8k 98.45 90.96 85.96 88.18 89.00
commonpool_l_laion_s1b_b8k 97.59 79.24 65.72 77.27 72.90
commonpool_l_s1b_b8k 98.02 88.72 81.65 86.36 87.20
commonpool_l_text_s1b_b8k 98.36 91.82 85.10 85.45 90.00
datacomp_l_s1b_b8k 98.19 80.10 69.51 71.82 76.60
datacomp_xl_s13b_b90k 99.14 68.39 53.14 62.73 66.80
laion2b_s34b_b88k 98.71 69.16 52.28 64.55 64.20
laion400m_e31 98.62 69.16 50.90 66.36 62.60
laion400m_e32 98.71 68.91 50.90 65.45 61.90
metaclip_400m 98.62 58.66 35.40 50.00 57.10
metaclip_fullcc 98.88 60.81 39.62 49.09 59.20
openai 97.59 39.71 20.24 33.64 37.60
ViT-B-16-SigLIP webli 99.22 81.40 75.45 82.73 78.10
ViT-B-16-SigLIP-256 webli 99.31 80.96 75.97 83.64 80.00
ViT-B-16-SigLIP-384 webli 99.48 84.24 80.45 84.55 82.20
ViT-B-16-SigLIP-512 webli 99.66 85.70 81.14 83.64 82.90
ViT-B-16-SigLIP-i18n-256 webli 99.40 89.66 85.70 89.09 85.70
ViT-B-16-plus-240 laion400m_e31 98.97 71.75 54.44 63.64 62.70
laion400m_e32 98.97 71.75 54.44 63.64 62.70
ViT-B-32 commonpool_m_basic_s128m_b4k 89.15 90.78 91.47 92.73 92.70
commonpool_m_clip_s128m_b4k 91.30 90.70 90.53 94.55 92.20
commonpool_m_image_s128m_b4k 89.66 91.30 92.08 94.55 92.40
commonpool_m_laion_s128m_b4k 89.23 90.35 90.18 94.55 91.20
commonpool_m_s128m_b4k 88.11 89.84 90.70 95.45 90.60
commonpool_m_text_s128m_b4k 89.32 91.04 90.96 94.55 93.40
commonpool_s_basic_s13m_b4k 72.70 74.33 74.76 72.73 74.10
commonpool_s_clip_s13m_b4k 76.31 76.57 76.66 76.36 77.10
commonpool_s_image_s13m_b4k 72.95 73.21 73.47 78.18 72.50
commonpool_s_laion_s13m_b4k 73.73 73.39 74.25 78.18 74.00
commonpool_s_s13m_b4k 74.68 74.25 73.99 70.91 72.40
commonpool_s_text_s13m_b4k 74.42 73.99 74.50 79.09 74.50
datacomp_m_s128m_b4k 89.75 90.61 90.61 94.55 92.20
datacomp_xl_s13b_b90k 98.88 79.33 64.17 69.09 76.30
laion2b_s34b_b79k 98.45 74.68 56.16 68.18 66.00
laion400m_e31 96.64 63.57 40.48 55.45 57.70
laion400m_e32 96.64 63.48 40.57 54.55 57.40
metaclip_400m 97.16 77.00 55.99 68.18 77.10
metaclip_fullcc 98.45 79.93 58.91 67.27 75.30
openai 96.47 60.81 33.59 52.73 51.40
ViT-B-32-256 datacomp_s34b_b86k 98.54 73.30 60.72 70.91 70.00
ViT-H-14 laion2b_s32b_b79k 99.22 62.96 47.63 51.82 55.60
metaclip_altogether 99.48 66.58 52.89 61.82 58.30
metaclip_fullcc 99.40 62.27 43.41 53.64 55.60
ViT-H-14-CLIPA datacomp1b 99.66 64.69 57.02 49.09 57.30
ViT-H-14-CLIPA-336 datacomp1b 99.66 61.15 55.64 46.36 59.10
laion2b 99.22 71.92 60.47 60.00 61.60
ViT-L-14 commonpool_xl_clip_s13b_b90k 99.48 57.62 48.84 44.55 51.80
commonpool_xl_laion_s13b_b90k 99.40 74.94 59.09 62.73 67.40
commonpool_xl_s13b_b90k 99.48 74.68 68.56 56.36 65.80
datacomp_xl_s13b_b90k 99.48 60.38 52.97 50.00 56.00
laion2b_s32b_b82k 99.05 62.53 48.23 59.09 56.40
laion400m_e31 99.05 67.18 51.42 63.64 59.50
laion400m_e32 99.05 66.75 52.02 63.64 59.90
metaclip_400m 98.71 57.71 39.62 50.00 54.30
metaclip_fullcc 99.14 63.14 45.48 47.27 52.60
openai 99.14 40.14 26.61 40.00 36.90
ViT-L-14-336 openai 99.22 33.85 26.44 33.64 39.90
ViT-L-14-CLIPA datacomp1b 99.48 72.61 63.57 63.64 66.80
ViT-L-14-CLIPA-336 datacomp1b 99.57 74.76 65.81 63.64 68.50
ViT-L-16-SigLIP-256 webli 99.57 86.48 80.10 81.82 79.60
ViT-L-16-SigLIP-384 webli 99.66 87.25 81.65 76.36 77.40
ViT-SO400M-14-SigLIP webli 99.74 84.07 82.17 84.55 81.10
ViT-SO400M-14-SigLIP-378 webli 99.66 85.62 85.01 83.64 81.30
ViT-SO400M-14-SigLIP-384 webli 99.66 85.44 84.58 81.82 81.20
ViT-SO400M-16-SigLIP-i18n-256 webli 99.74 88.89 83.98 85.45 86.30
ViT-bigG-14 laion2b_s39b_b160k 99.40 70.89 58.91 56.36 58.50
metaclip_fullcc 99.31 68.04 53.66 60.00 58.30
ViT-bigG-14-CLIPA datacomp1b 99.66 68.04 59.35 56.36 63.20
ViT-bigG-14-CLIPA-336 datacomp1b 99.74 64.00 56.68 58.18 59.40
ViT-g-14 laion2b_s12b_b42k 99.14 69.25 55.38 59.09 54.70
laion2b_s34b_b88k 99.05 61.93 47.03 54.55 50.10
convnext_base laion400m_s13b_b51k 98.54 59.09 40.40 49.09 54.10
convnext_base_w laion2b_s13b_b82k 98.79 66.84 52.28 60.91 59.40
laion2b_s13b_b82k_augreg 98.88 68.73 56.33 60.00 57.40
laion_aesthetic_s13b_b82k 98.45 68.13 51.59 59.09 59.70
convnext_base_w_320 laion_aesthetic_s13b_b82k 98.62 68.91 51.25 55.45 61.40
laion_aesthetic_s13b_b82k_augreg 98.62 72.27 59.00 59.09 63.90
convnext_large_d laion2b_s26b_b102k_augreg 99.40 67.18 54.78 54.55 53.70
convnext_large_d_320 laion2b_s29b_b131k_ft 99.31 67.10 57.11 55.45 56.90
laion2b_s29b_b131k_ft_soup 99.57 67.53 56.16 53.64 57.00
convnext_xxlarge laion2b_s34b_b82k_augreg 99.48 64.94 54.69 50.91 59.80
laion2b_s34b_b82k_augreg_rewind 99.22 66.15 54.09 55.45 58.80
laion2b_s34b_b82k_augreg_soup 99.48 64.94 54.44 55.45 59.00
nllb-clip-base v1 94.66 76.66 61.67 71.82 74.90
nllb-clip-base-siglip mrl 98.28 71.92 63.91 70.91 69.40
v1 98.71 77.69 75.88 75.45 78.60
nllb-clip-large v1 97.50 61.41 48.58 56.36 60.90
nllb-clip-large-siglip mrl 98.88 72.27 66.75 66.36 71.20
v1 99.05 73.90 75.37 74.55 75.90
roberta-ViT-B-32 laion2b_s12b_b32k 97.67 75.45 57.45 62.73 68.50
xlm-roberta-base-ViT-B-32 laion5b_s13b_b90k 98.62 76.83 59.00 69.09 71.20
xlm-roberta-large-ViT-H-14 frozen_laion5b_s13b_b90k 99.22 63.74 47.89 54.55 56.70
llava-llama3-1.1:8b - 98.09 39.50 44.04 48.62 41.83
llava-1.6:7b - 97.50 57.25 61.45 53.64 51.20
llava-1.6:13b - 98.88 58.00 55.42 54.55 56.00
llava-1.6:34b - 98.97 84.85 85.11 89.09 81.50
llama3.2-vision:90b - 98.88 71.01 74.65 77.06 68.41
gpt-4o-mini-2024-07-18 - 99.40 84.68 87.09 90.91 85.15
gpt-4o-2024-08-06 - 99.48 96.82 94.58 99.09 96.73